# Parametric tuning of dynamical phase transitions in ultracold reactions

### v1

@@ Abstract
Advances in ultracold chemistry have led to the possibility of a coherent transformation between ultracold atoms and molecules including between completely bosonic condensates. Such transformations are enabled by the magneto-association of atoms at a Feshbach resonance which results in a passage through a quantum critical point. In this study, we show that the presence of generic interaction between the formed molecules can fundamentally alter the nature of the critical point, change the yield of the reaction and the order of the consequent phase transition. We find that the correlations introduced by this rather general interaction induce nontrivial many-body physics such as coherent oscillations between atoms and molecules, and a selective formation of squeezed molecular quantum states and quantum cat states. We provide analytical and numerical descriptions of these many-body effects, along with scaling laws for the reaction yield in both the adiabatic and non-adiabatic regimes, and highlight the potential experimental relevance in quantum sensing.
preprint: APS/123-QED
@@ I Introduction
Ultracold chemistry holds a prominent place in quantum engineering [1], information processing [2], and further understanding of fundamental quantum phenomena such as superconductivity, magnetism and mechanisms of elementary chemical reactions. One of the landmark experiments in the field of ultracold chemistry is the creation of Bose-Einstein condensates (BECs) using dilute gases of alkali metal atoms [3, 4, 5] in the 1990s. A series of experimental developments followed in the next two decades which led to major strides of progress in ultracold chemistry, quantum control and finally in the creation of a degenerate Fermi gas [6] and degenerate Bose gas of molecules [7], and observation of Bosonic stimulation of this reactive transformation [8].
Central to almost all the experimental developments concerning the creation of ultracold molecules is the concept of Feshbach resonance that allows for the association of bosonic/fermionic atoms by using a tunable magnetic field. Tuning a magnetic field across the Feshbach resonance drives an ultracold atomic system through a quantum critical point and hence a dynamical phase transition [9, 10] which results in the conversion of atoms to molecules. This ‘reaction’ is quantum coherent and reversible at macroscopic scale. The time-dependent Tavis-Cummings (TC) model [11] is a minimal model that captures many nontrivial many-body effects that are encountered during the stimulated reaction. Its Hamiltonian is
@@
Here, $S$ ($\gg 1$) is the quantum mechanical pseudo-spin, and $\beta$ is the sweep rate of the transition through the Feshbach resonance. The sweep rate is controllable experimentally in a broad range – from almost instantaneous to quasi-adiabatic. The bosonic operator, $\hat{\psi}^{\dagger}$, creates a molecule. The terms with the pseudo-spin raising and lowering operators, $\hat{S}^{+}$ and $\hat{S}^{-}$, correspond, respectively, to the dissociation of a molecule and association of two fermionic atoms to form a bosonic molecule with a characteristic coupling $g$ (see [12, 13] for derivation of this model from true atomic-molecular Hamiltonians). The driven TC model can also be reformulated in terms of a fully bosonic reaction between bosonic atomic and molecular condensates [12, 13]. We will focus here on the reaction in which there are no molecules initially. This corresponds in (1) to the initial state, as $t\rightarrow-\infty$, without molecules and the spin fully up-polarized along the $z$-axis.
@@
Figure 1: An ultracold atomic reaction undergoing, respectively, (a),(b) the first-order and (d),(e) the second-order dynamical phase transition. (a) For the first-order transition, both molecules and atoms can be in the local energy minima simultaneously, whereas (d), for the second-order transition, a high efficiency of the chemical reaction can be achieved by an adiabatic transition between the molecular and atomic energy minima. (b) and (e) The mean-field ground energy as a function of the order parameter in the first-order and second-order dynamical phase transition, respectively (arrows indicate the direction of time) (c) The numerically obtained phase diagram describing the dependence of the number of nonadiabatic excitations, $n_{\text{ex}}$, as a function of the inverse sweep rate, $1/\beta_{\rm eff}\equiv g^{2}/(\beta\log_{e}N)$ ($x$-axis), and the molecular interaction strength, $\kappa=r/g$ in Eqs. (1) and (3), ($y$-axis). For the model (2), $\kappa=-1$ separates the regimes with first ($\kappa<-1$) and second ($\kappa>-1$) order phase transitions. This explains the fast increase of the excitation numbers below the $\kappa=-1$ line.
The time-dependent sweep is needed in practice to make all atoms encounter the resonance. In addition, the quantum adiabatic theorem guarantees that a sufficiently slow sweep converts the initial atomic state into the bosonic ground state, which is the molecular condensate. Hence, potentially, a 100% efficiency of the reaction is experimentally possible. However, the true adiabatic limit cannot be reached, so it is important to understand the quasiadiabatic regime with small but finite $\beta$. This regime is characterized by the number of nonadiabatic excitations, $n_{ex}$, which is the number of the molecules that would appear in the adiabatic limit but did not form after the transition through the resonance, i.e., as $t\rightarrow+\infty$.
An unusual theoretical finding about the time-dependent TC model was the discovery of its integrability [14, 15], which proved the existence of a dynamical phase transition and confirmed semiclassical predictions for the power-law scaling of the nonadiabatic excitation density after the quasi-adiabatic sweep through the resonance [12]. However, recent experiments [8] with a bosonic reaction show a behavior that is not known within the slowly driven TC model, such as coherent oscillations between atoms and molecules after crossing the Feshbach resonance. Moreover, even with the slowest sweep rates, a system usually ends up in a pre-thermalized atomic-molecular mixture state, in disagreement with the nearly perfect reaction efficiency predicted by the TC model [16].
One possibility to extend the TC model to account for the richer reaction dynamics is to add a dispersion of the atomic states. This results in a generalized TC model [15, 13] that, surprisingly, is also solvable but leads to essentially the same predictions for the behavior of the nonadiabatic excitations on the sweep rate as the minimal model (1). Hence, other interaction types may be responsible for the experimentally observed behavior beyond the standard TC model.
In this article, we consider a different generalization of the model (1), in which we add another molecular coupling term:
@@
and where $\hat{n}$ is the molecular number operator; $r$ is the coupling strength. The coupling in (3) is effectively broadening the molecular dispersion energy, which becomes now dependent on the number of molecules that changes during the reaction. Such a coupling must be present in ultracold molecules due to dipole interactions and elastic scatterings.
@@
We find that this addition to the model leads to a substantially richer behavior, especially for quasi-adiabatic transitions. This could be anticipated from the applications of the TC model in optics, where such a nonlinear term describes the optical mode in a ‘Kerr-like’ medium [17, 18, 19].
@@
With this addition, the phase diagram for the reaction efficiency depends on both the sweep rate $\beta$ and the nonlinearity $r$. In Fig. 1 we show the result of our numerical simulations for the number of the nonadiabatic excitations, which is the number of unformed molecules, following the sweep of the chemical potential. By setting $r=\kappa g$, the line $\kappa=-1$ marks a critical nonlinearity, such that for $\kappa<-1$ the number of the nonadiabatic excitations is much larger than for $\kappa>-1$, especially in the quasi-adiabatic limit (large $1/\beta_{\rm eff}\equiv g^{2}/(\beta\log_{e}N)$). We will show that this behavior follows from the possibility of either a second- or first-order phase transition during the sweep through the Feshbach resonance. Our model is no longer exactly solvable but in the quasi-adiabatic limit it can be studied analytically in detail.
II Semiclassical description and phase transitions
The experimentally relevant regime corresponds to a large number of molecules, $N=2S\gg 1$, that can be potentially created during the sweep through the resonance. This justifies the semiclassical approximation. To develop it, first, we note that our model conserves
@@ II Semiclassical description and phase transitions
whose eigenvalue $N=2S$ corresponds to our initial conditions, where the starting state |$\ket{\psi_{0}}$ is the one without molecules. Hence, it is convenient to mark all states by the number of molecules:
@@
where $|n\rangle_{m}$ is the state with $n$ molecules and $|S-n\rangle_{S_{z}}$ is the spin state with $S_{z}=S-n$. The initial state as $t\rightarrow-\infty$ corresponds to $|n\rangle=|0\rangle$. The matrix elements $\langle n|\hat{H}_{TC}+\hat{H}_{\rm int}|m\rangle$ in this basis are given by
$\begin{split}H_{nm}=(-\beta tn+\frac{r}{N}n^{2})\>\delta_{n,m}+gn\sqrt{\frac{N-n+1}{N}}\>\delta_{n,m-1}\\ +\;g(n+1)\sqrt{\frac{N-n}{N}}\>\delta_{n,m+1}.\end{split}$ (5)
Let us look for the solution to the Schrödinger equation of the form $|\psi\rangle=\sum_{n=0}^{N}a_{n}(t)|n\rangle$, and introduce the amplitude generating function
$u(\phi,t)=\sum_{n=0}^{N}a_{n}(t)e^{in\phi}.$ (6)
Note that $na_{n}e^{in\phi}=-ia_{n}\partial e^{in\phi}/\partial\phi$, so the time-dependent Schrödinger equation for the amplitudes can be written in terms of a single equation for $u(\phi,t)$ as
@@
where we associate $\hat{n}\equiv-i\partial/\partial\phi$. In the semiclassical approximation we can then associate $\phi$ with a coordinate that is conjugate to the classical momentum $n$. In addition, we disregard the terms of order $1/N$. Then, the classical Hamiltonian that corresponds to the Schrödinger equation (7) has the form
$H_{cl}(n,\phi;\gamma)=\gamma n+\frac{r}{N}n^{2}+2gn\sqrt{\frac{N-n}{N}}\cos(\phi),$ (8)
with a time-dependent parameter
$\gamma=\gamma(t)=-\beta t,$
and the classical equations of motion
@@
III Quasi-adiabatic second-order phase transition
Up to the new $r$-dependent nonlinear term, the Hamiltonian (8) coincides with the analogous semiclassical Hamiltonian in [12]. The relation between the classical variables and the number of the nonadiabatic excitations is established by noting that as $t\rightarrow\pm\infty$ the time-dependent term completely dominates: $H_{cl}\sim\gamma n$. Following [12], we note that the equations of motion initially conserve $n$ and the adiabatic invariant is given by the initial number of molecules, $n_{-\infty}$:
@@ III Quasi-adiabatic second-order phase transition
If during the evolution the adiabatic invariant acquires a small contribution $\Delta I$, this is interpreted in the semiclassical theory as the density of the produced elementary nonadiabatic excitations, i.e, $n_{ex}=\Delta I/N$. In the strict adiabatic limit, the ground state with no molecules eventually is transferred into the new ground state with $N$ molecules. Taking into account the nonadiabatic excitations, the number of the created molecules in our process is given by
@@
We should assume that initially $n_{-\infty}=I_{-\infty}\sim 1$, which is negligible in comparison to the large $N$. Next, we assume a nearly adiabatic sweep of the chemical potential [12]. The point $\phi^{*}=\pi$ is a steady point of the classical equations (9), in the vicinity of which the system evolves. Assuming that near this point $n\ll N$, and retaining the terms of the lowest order we find an effective Hamiltonian that governs the evolution at the early stage of the reaction, that is
$H_{cl}(n,\phi;\gamma)\approx\left(\gamma-2g\right)n+\left(\frac{r+g}{N}\right)n^{2}+gn\phi^{2}.$ (10)
For a quasi-adiabatic evolution, the nonadiabatic excitations are generally suppressed exponentially, as $\Delta I\sim\exp(-a/\beta)$, with some finite positive $a$. Such excitations are not essential and we can safely disregard them. However, according to the Kibble-Zurek phenomenology, the excitations are enhanced near a critical point, at which the symmetry of the original ground state breaks down spontaneously. It turns out that the Hamiltonian (10) contains this event, so it is sufficient to describe the phase transition quantitatively.
@@
The initial conditions correspond to $Q\sim 1$. In fact, a quantum mechanical treatment of the initial conditions needs averaging of the behavior over a distribution of small initial values with $0<Q^{2}\ll N$ [12]. However, we will show later that this information is irrelevant within the leading order semiclassical description. The system with the potential (14) experiences the 2nd order phase transition at $s=0$. Indeed, for $s<0$, the potential $V(Q)$ has a single minimum at $Q=0$ but for $s>0$, there are two local minima at $Q_{min}\sim\pm\sqrt{\beta Ns/[2(r+g)]}$.
In Fig. 1, we illustrate that for $s<0$ the system is initially near a single minimum but it follows one of the new minima for $s>0$. In phase space, this corresponds to crossing a separatrix, in which vicinity the classical adiabatic invariant is no longer conserved. Thus, so far our approximations were justified because they capture the main source of the nonadiabatic excitations near the phase transition.
The evolution equations for $Q$ and $P$ with the Hamiltonian (10), acquire a universal form after re-scaling of the variables:
@@
The rescaling of variables, however, was not canonical, so it did not conserve the action:
@@
Hence, if $I$ is the adiabatic invariant in the original variables, $n$ and $\phi$, then in the rescaled $Q$ and $P$ the invariant is given by ${\cal I}=I/(uv)$. For our case, we found
@@
$\Delta I_{r=0}=\frac{uv}{4\pi}\log_{e}\left(\frac{uv}{4I_{-}\pi}\right)=\frac{N\beta}{2\pi g^{2}}\log_{e}\left(\frac{N\beta}{2\pi g^{2}I_{-}}\right),$ (18)
where either the exact solution or further semiclassical analysis can be used to fix $I_{-}\approx 1$ for our initial conditions in the model (1). According to [15], this result is valid only in the quasi-adiabatic case, namely for $1/\log_{e}N\gg\beta/(2\pi g^{2})\gg 1/N$. For example, at $I_{-}=1$, in this range Eq. (18) coincides with the exactly found expression for the position of the maximum of the probability distribution of the excitation number.
@@
$\Delta I(r)=\frac{N\beta}{2\pi g(g+r)}\log_{e}\left(\frac{N\beta}{2\pi g(g+r)I_{-}}\right),$ (19)
where $I_{-}$ is the parameter of order $1$ that characterizes the initial conditions. As this unknown factor appears inside the logarithm in (19), its precise value is not important because its relative contribution to (19) vanishes in the limit $N\rightarrow\infty$ for the quasi-adiabatic evolution with $1/\log_{e}N\gg\beta/(2\pi(g^{2}+rg))\gg 1/N$. For comparison with numerical results, we set $I_{-}=1$.

### v2

# Parametric tuning of quantum phase transitions in ultracold reactions

@@ Abstract
Advances in atomic physics have led to the possibility of a coherent transformation between ultracold atoms and molecules including between completely bosonic condensates. Such transformations are enabled by the magneto-association of atoms at a Feshbach resonance which results in a passage through a quantum critical point. In this study, we show that the presence of generic interaction between the constituent atoms and molecules can fundamentally alter the nature of the critical point, change the yield of the reaction and the order of the consequent phase transition. We find that the correlations introduced by this interaction induce nontrivial many-body physics such as coherent oscillations between atoms and molecules, and a selective formation of squeezed molecular quantum states and quantum cat states. We provide analytical and numerical descriptions of these effects, along with scaling laws for the reaction yield in non-adiabatic regimes.
@@ I Introduction
Ultracold atoms hold a prominent place in quantum engineering [1], information processing [2], and further understanding of fundamental quantum phenomena such as superconductivity, magnetism and mechanisms of elementary chemical reactions. One of the landmark experiments in the field is the creation of Bose-Einstein condensates (BECs) using dilute gases of alkali metal atoms [3, 4, 5] in the 1990s. A series of developments have led to the creation of a degenerate Fermi gas [6]. More recently, it became possible to observe a quantum coherent chemical reaction for a macroscopic number of bosonic atoms forming a degenerate Bose gas of molecules [7, 8].
The association of bosonic/fermionic atoms into molecules is achieved by tuning a time-dependent magnetic field across the Feshbach resonance, which drives an ultracold atomic system through a quantum phase transition [9, 10]. This ‘reaction’ is quantum coherent and reversible at macroscopic scale. The time-dependent Tavis-Cummings (TC) model [11] is a minimal model that captures many non-trivial many-body effects that are encountered during the stimulated reaction. Its Hamiltonian is
@@
Here, $S$ ($\gg 1$) is the quantum mechanical pseudo-spin, and $\beta$ is the sweep rate of the transition through the Feshbach resonance. The sweep rate is experimentally controllable in a broad range – from almost instantaneous to quasi-adiabatic. The bosonic operator, $\hat{\psi}^{\dagger}$, creates a molecule. The terms with the pseudo-spin raising and lowering operators, $\hat{S}^{+}$ and $\hat{S}^{-}$, correspond, respectively, to the dissociation of a molecule and association of two fermionic atoms to form a bosonic molecule with a characteristic coupling constant $g$. The driven TC model can also be reformulated in terms of a fully bosonic reaction between bosonic atomic and molecular condensate (see [12, 13] and section IV.1 for derivation of this model from true atomic-molecular Hamiltonians). We will focus here on the reaction in which there are no molecules initially. This corresponds in (1) to the initial state, as $t\rightarrow-\infty$, without molecules and the spin fully up-polarized along the $z$-axis.
The time-dependent field sweep is needed in practice to make all atoms encounter the resonance. In addition, the quantum adiabatic theorem guarantees that a sufficiently slow sweep converts the initial atomic state into the bosonic ground state, which is the molecular condensate. Hence, potentially, a 100% efficiency of the reaction is experimentally possible. However, the true adiabatic limit cannot be reached, so it is important to understand the quasiadiabatic regime with a small but finite $\beta$. This regime is characterized by the number of nonadiabatic excitations, $n_{ex}$, which is the number of the molecules that would appear in the adiabatic limit but did not form after the transition through the resonance, i.e., as $t\rightarrow+\infty$.
An unusual theoretical finding about the time-dependent TC model was the discovery of its integrability [14, 15], which confirmed semiclassical predictions for the power-law scaling of the nonadiabatic excitation density after the quasi-adiabatic sweep through the resonance [12].
However, recent experiments [8] with a bosonic reaction show a behavior that is not known within the slowly driven TC model, such as coherent oscillations between atoms and molecules after crossing the Feshbach resonance. Moreover, even with the slowest sweep rates, a system usually ends up in a prethermalized atomic-molecular mixture state, in disagreement with the nearly perfect reaction efficiency predicted by the TC model [16].
One possibility to extend the TC model to account for the richer reaction dynamics is to add a dispersion of the atomic states. This results in a generalized TC model [15, 13] that, surprisingly, is also solvable and leads to essentially the same predictions for the dependence of the number of nonadiabatic excitations on the sweep rate as the minimal model (1). Hence, other interaction types may be responsible for the experimentally observed behavior beyond the standard TC model.
@@
Figure 1: An ultracold atomic reaction undergoing a (non-adiabatic/quasi-adiabatic) passage through, respectively, (a),(b) the first-order and (d),(e) the second-order quantum phase transition. (a) For the first-order transition, both molecules and atoms can be in the local energy minima simultaneously, whereas (d), for the second-order transition, a high efficiency of the chemical reaction can be achieved by an adiabatic transition between the molecular and atomic energy minima. (b) and (e) The mean-field ground energy as a function of the order parameter in the first-order and second-order quantum phase transition, respectively (arrows indicate the direction of time) (c) The numerically obtained phase diagram describing the dependence of the number of nonadiabatic excitations, $n_{\text{ex}}$, as a function of the inverse sweep rate, $1/\beta_{\rm eff}\equiv g^{2}/(\beta\ln N)$ ($x$-axis), and the molecular interaction strength, $\kappa=r/g$ in Eqs. (1) and (3), ($y$-axis). For the model (2), $\kappa=-1$ separates the regimes with first ($\kappa<-1$) and second ($\kappa>-1$) order phase transitions. This explains the fast increase of the excitation numbers below the $\kappa=-1$ line.
In this article, we consider a different generalization of the model (1), in which we add another interaction term:
@@
and where $\hat{n}$ is the molecular number operator; $r$ is the interaction strength. This interaction term includes all the ‘non-resonant’ interactions that are possible due to elastic scattering between atoms and molecules and is hence expected to have the same level of significance as the resonance term in (1). Specifically, this interaction term consists of two-body atom-atom, atom-molecule and molecule-molecule scattering terms, all of which can be accounted for using the molecule number operator $\hat{n}$ due to the fact that quantum dynamics conserves the number of particles (explained in detail in sections II.1 and IV.2) . The coupling in (3) effectively broadens the molecular dispersion energy, which becomes now dependent on the number of molecules that changes during the reaction. Such a coupling must be present in ultracold molecules due to elastic scattering and dipole interactions, which in certain atomic systems can be stronger than the coupling induced by the Feshbach resonance [17, 18].
The main reason for the inclusion of the interaction term in (3) is the fact that the TC model (1) is a phenomenological model that was introduced as a simplification of a system of bosonic atoms in a confinement with complex pairwise atomic interactions. The full many-body Hamiltonian includes only the interacting atoms in a trap. To obtain the TC model from that Hamiltonian, one should identify a quasi-stable bound state (the molecule), and then keep only the interactions of this bound state with free atomic states (see sections IV.1 and IV.2). This truncation of the original Hamiltonian is justified not by the dominance of the coupling $g$ over all other interactions but rather by the fact that the creation of the molecules from the atoms is a distinct effect, whereas the remaining interactions, the most relevant of which is given by Eq. (3), are usually assumed to just renormalize the parameters of quasiparticles in a condensed matter.
@@
We will show, however, that such an omission is not justified when considering a transition through a critical point, particularly when the characteristic energy scales for $r$ and $g$ are comparable. The addition of $\hat{H}_{\rm int}$ leads to substantially richer behavior, especially for quasi-adiabatic transitions. This could be anticipated from the applications of the TC model in optics, where such a nonlinear term describes the optical mode in a ‘Kerr-like’ medium [19, 20, 21].
@@
With this addition, the phase diagram for the reaction efficiency depends on both the sweep rate $\beta$ and the nonlinearity $r$. In Fig. 1 we show the result of our numerical simulations for the number of the nonadiabatic excitations, which is the number of unformed molecules, following the sweep of the chemical potential. We set $r=\kappa g$, where the line $\kappa=-1$ marks a critical nonlinearity; for $\kappa<-1$ the number of the nonadiabatic excitations is much larger than for $\kappa>-1$, especially in the quasi-adiabatic limit (large $1/\beta_{\rm eff}\equiv g^{2}/(\beta\log_{e}N)$). We will show that this behavior follows from the possibility of either a second- or first-order quantum phase transition during the sweep through the Feshbach resonance. Our model is no longer exactly solvable but in the quasi-adiabatic limit it can be studied analytically in detail. We propose our model as a guide for future experiments that explore the complex quantum dynamics of interacting molecular condensates. In particular, we will derive the formulas for the efficiency of the reaction between molecular and atomic condensates as a function of the sweep rate $\beta$, which can be also used to estimate the parameters $r$ and $g$ experimentally (see also  [22, 23, 24, 25, 26, 27]). We also show that our approach can be extended to more complex types of Feshbach resonance, such as the one considered in the experiment described in [8].
II Results
The experimentally relevant regime corresponds to a large number of starting atoms, $N=2S\gg 1$, that can potentially create a large number of molecules during the sweep through the resonance. This forms the basis of our semiclassical approximation, which we develop as follows. First, we rewrite the Schrödinger equation for the nonlinear TC model (2) in the form of the Schrödinger equation of a particle in a one-dimensional potential. The large $N$-limit can then be understood by considering the classical equations of motion in the same potential. Next we show that the potential energy changes with time: for the parameter regime $r+g>0$, it starts off as a harmonic oscillator potential as $t\rightarrow-\infty$, which is converted into a double-well potential during the passage through a second-order phase transition. The number of the nonadiabatic excitations generated during this passage is obtained from the change of the adiabatic invariant upon this transition.
This step reveals the universality of the equations of motion near the critical point for $r+g>0$, so we extend the known exact formula for the TC model at $r=0$ to the case with $r\neq 0$. We then show that for sufficiently strong attractive interactions (i.e. $r+g<0$), there is a qualitative change in the behavior of the nonlinear TC model, due to the emergence of an underlying first-order phase transition. In this case, the potential energy changes from a ‘triple well’ potential before the transition to a double well potential after the transition. We derive scaling relations for the number of nonadiabatic excitations which do not vanish even in the adiabatic limit, which is a characteristic of the first-order transition. We then demonstrate some of the dynamical features accompanying this transition, such as oscillatory dynamics of the excitations and formation of nonclassical states, using numerical simulations.
II.1 Semiclassical description and phase transitions
Note that our model conserves
@@ II.1 Semiclassical description and phase transitions
whose eigenvalue $N=2S$ corresponds to our initial conditions, where the starting state is the one without molecules. Hence, it is convenient to mark all states by the number of molecules:
@@
where $|n\rangle_{m}$ is the Fock state with $n$ bosonic molecules and $|S-n\rangle_{S}$ is the pseudospin state with z-axis projection $S_{z}=S-n$. For example, $\hat{\psi}|n\rangle_{m}=\sqrt{n}|n-1\rangle_{m}$, $S_{z}|S-n\rangle_{S}=(S-n)|S-n\rangle_{S}$, $S^{+}|S_{z}\rangle_{S}=\sqrt{S(S+1)-S_{z}(S_{z}+1)}|S_{z}+1\rangle_{S}$ and so on. Note also that due to the conservation law (4), the addition of the interatomic interactions of the form $\sim S_{z}^{2}$ and atom-molecule (non-resonant) interactions of the form $nS_{z}$ to the original Hamiltonian would be equivalent to a redefinition of the coupling $r$ in the term (3) and a time shift (see section IV.2). We assume that this redefinition is already included in our definition of $r$.
The initial state as $t\rightarrow-\infty$ corresponds to $|n\rangle=|0\rangle$. The matrix elements $H_{nm}\equiv\langle n|\hat{H}_{TC}+\hat{H}_{\rm int}|m\rangle$ in this basis are given by
$\begin{split}H_{nm}=&\>(-\beta tn+\frac{r}{N}n^{2})\>\delta_{n,m}+gn\sqrt{\frac{N-n+1}{N}}\>\delta_{n,m-1}\\ &+g(n+1)\sqrt{\frac{N-n}{N}}\>\delta_{n,m+1}\end{split}$ (5)
We will show that the critical point appears at conditions $1\ll n\ll N$, so by disregarding terms of order $(n/N)^{3}$ and $1/N$, the matrix elements can be written as
$\begin{split}H_{nm}=&\>(-\beta tn+\frac{r}{N}n^{2})\>\delta_{n,m}+\\ &gn\left(1-\frac{n}{2N}\right)(\delta_{n,m-1}+\delta_{n,m+1})+O(1/n).\end{split}$ (6)
A formal solution to the time-dependent Schrödinger equation for the driven nonlinear TC model (2) can be written as: $|\psi\rangle=\sum_{n=0}^{N}a_{n}(t)|n\rangle$. We introduce the amplitude generating function
$u(\phi,t)=\sum_{n=0}^{N}a_{n}(t)e^{in\phi},$ (7)
whose name follows from the fact that the amplitudes can be obtained by the inverse Fourier transformation
[…]
For the Hamiltonian (6), the time-evolution of the amplitudes, up to the omitted terms $\sim 1/n$, is given by
$i\dot{a}_{n}=(-\beta tn+\frac{r}{N}n^{2})a_{n}+gn\left(1-\frac{n}{2N}\right)(a_{n-1}+a_{n}).$ (8)
Note that $na_{n}e^{in\phi}=-ia_{n}\partial e^{in\phi}/\partial\phi$. We then multiply Eq. (8) by $e^{in\phi}$ and sum all equations for $n=0,1,2,\ldots$. Using the definition in (7), the time-dependent Schrödinger equation for the amplitudes is then written in terms of a single equation for $u(\phi,t)$ as
@@
where we associate $\hat{n}\equiv-i\partial/\partial\phi$. In the semiclassical approximation we can then associate $\phi$ with a coordinate that is conjugate to the classical momentum $n$. Then, the classical Hamiltonian that corresponds to the Schrödinger equation (9) has the form
$H_{cl}(n,\phi)=-\beta tn+\frac{r}{N}n^{2}+2gn\left(1-\frac{n}{2N}\right)\cos(\phi),$ (10)
and the classical equations of motion are given by
@@
II.2 Quasi-adiabatic second-order phase transition
Without the new $r$-dependent nonlinear term, the Hamiltonian (10) would coincide with the analogous semiclassical Hamiltonian in [12]. The relation between the classical variables and the number of nonadiabatic excitations is established by noting that as $t\rightarrow\pm\infty$ the time-dependent term completely dominates: $H_{cl}\sim-\beta tn$. Following [12], we note that the equations of motion initially conserve $n$ and the adiabatic invariant is given by the initial number of molecules, $n_{-\infty}$:
@@ II.2 Quasi-adiabatic second-order phase transition
If during the evolution the adiabatic invariant acquires a small contribution $\Delta I$, this is interpreted in the semiclassical theory as the density of the produced elementary nonadiabatic excitations, i.e, $n_{ex}=\Delta I/N$. In the strict adiabatic limit, the ground state with no molecules eventually transitions into the new ground state with $N$ molecules. Taking the nonadiabatic excitations into account, the number of molecules created in our process is given by
@@
We should assume that initially $n_{-\infty}=I_{-\infty}\sim 1$, which is negligible in comparison to the large $N$. We assume a nearly adiabatic sweep of the chemical potential [12]. The point $\phi^{*}=\pi$ is a steady point of the classical equations (11), in the vicinity of which the system evolves. Assuming that near this point $n\ll N$, and retaining the terms of the lowest order we find an effective Hamiltonian that governs the evolution at the early stage of the reaction, that is
$H_{cl}(n,\phi;\gamma)\approx\left(-\beta t-2g\right)n+\left(\frac{r+g}{N}\right)n^{2}+gn\phi^{2}.$ (12)
For a quasi-adiabatic evolution, the nonadiabatic excitations are generally suppressed exponentially, as $\Delta I\sim\exp(-a/\beta)$, with some finite positive $a$. Such excitations are not essential and we can safely disregard them. However, according to the Kibble-Zurek phenomenology, the excitations are enhanced near a critical point at which the symmetry of the original ground state breaks down spontaneously. It turns out that the Hamiltonian (12) contains this event, so it is sufficient to describe the phase transition quantitatively.
@@
The initial conditions correspond to $Q\sim 1$. In fact, a quantum mechanical treatment of the initial conditions requires the averaging of the behavior over a distribution of small initial values with $0<Q^{2}\ll N$ [12]. However, we will show later that this information is irrelevant within the leading-order semiclassical description. The system with the potential (16) experiences the second-order phase transition at $s=0$. Indeed, for $s<0$, the potential $V(Q)$ has a single minimum at $Q=0$ but for $s>0$, there are two local minima at $Q_{min}\sim\pm\sqrt{\beta Ns/[2(r+g)]}$.
In Fig. 1, we illustrate that for $s<0$ the system is initially near a single minimum but it follows one of the new minima for $s>0$. In phase space, this corresponds to crossing a separatrix, in whose vicinity the classical adiabatic invariant is no longer conserved. Thus, so far our approximations were justified because they capture the main source of the nonadiabatic excitations near the phase transition.
The evolution equations for $Q$ and $P$ with the Hamiltonian (12), acquire a universal form after rescaling of the variables:
@@
The rescaling of variables, however, is not canonical, so it does not conserve the action:
@@
Hence, if $I$ is the adiabatic invariant in the original variables, $n$ and $\phi$, then in the rescaled $Q$ and $P$ coordinates, the invariant is given by ${\cal I}=I/(uv)$. For our case, we found
@@
$\Delta I_{r=0}=\frac{uv}{4\pi}\ln\left(\frac{uv}{4I_{-}\pi}\right)=\frac{N\beta}{2\pi g^{2}}\ln\left(\frac{N\beta}{2\pi g^{2}I_{-}}\right),$ (20)
where either the exact solution or further semiclassical analysis can be used to fix $I_{-}\approx 1$ for our initial conditions in the model (1). According to [15], this result is valid only in the quasi-adiabatic case, namely for $1/\ln N\gg\beta/(2\pi g^{2})\gg 1/N$. For example, at $I_{-}=1$, in this range Eq. (20) coincides with the exact expression for the position of the maximum of the probability distribution of the excitation number.
@@
$\Delta I(r)=\frac{N\beta}{2\pi g(g+r)}\ln\left(\frac{N\beta}{2\pi g(g+r)I_{-}}\right),$ (21)
where $I_{-}$ is the parameter of order $1$ that characterizes the initial conditions. As this unknown factor appears inside the logarithm in (21), its precise value is not important because its relative contribution to (21) vanishes in the limit $N\rightarrow\infty$ for the quasi-adiabatic evolution with $1/\ln N\gg\beta/(2\pi(g^{2}+rg))\gg 1/N$. For comparison with numerical results, we set $I_{-}=1$.